# Multi-Agent Collaboration Framework for Recommender Systems

### v1

@@ Abstract.
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYprice: 15.00isbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
Researchers have started to utilize the capabilities of agents to solve recommendation tasks. Existing work like  (Zhang et al., 2023c; Wang et al., 2023b; Zhang et al., 2023b) primarily focuses on employing agents for simulating user or item behaviors, providing insights into user preferences but falling short of integration into RSs. On the other hand, some studies (Wang et al., 2023a; Huang et al., 2023) attempt to leverage the capabilities of agents to directly build a recommender, primarily using one single agent with planning and memory components and auxiliary tools (e.g., search engine). However, there are various complex decision-making tasks in recommendation scenarios, on which single-agent instances are unable to perform well. Multi-agent collaboration, which is near to human workflows, is believed to accomplish complex tasks better with collective intelligence. Although work (Shu et al., 2023) proposes a multi-agent recommendation framework, it only has limited agent types and a fixed collaboration mode.
@@ 3.1. Framework Overview
The following sections will detail each agent’s specific characteristics and functions.
[The framework of MACRec.]The framework of MACRec.
@@ 3.2.1. Manager
For the given task, the Manager would assign sub-tasks to other agents and complete the main execution process. It oversees the coordination and collaboration among all other agents, acting like a manager.
@@ 3.2.2. Reflector
The Reflector will step in when the Manager is about to perform the second or more run on the same task input. If the Reflector judges that the answer given by the Manager has no room for improvement, the Manager will no longer perform the current run. Otherwise, the Reflector will further summarize where the Manager can be improved, e.g., not considering the few highly rated items/movies in the user’s historical interactions.
@@ 3.2.5. Task Interpreter
The Task Interpreter will get the history of the user’s conversations with the system when starts running. Since conversation histories can be long, the Task Interpreter will only get the last part of the history. The Task Interpreter also has access to call the text summarization tool to get a more concise overview of the history. Eventually, the Task Interpreter will give a specific description of the task requirements that will be used to guide the subsequent runs of the Manager.
@@ 4.2. Sequential Recommendation (SR)
Modeling of user’s long-term and short-term interests is important in sequential recommendation tasks. Hence, the User Analyst’s role is self-evident. The number of relevant items in the sequence recommendation is significantly higher than the rating prediction task. It is hard to ask the Item Analyst to analyze every item that appeared in either the history or the candidate set. Therefore, the role of the Item Analyst is comparatively limited compared to other scenarios. Moreover, given that the answers to the sequential recommendation task are much more complex (i.e., a ranking order of the candidate set), the Reflector can help to avoid the Manager getting into formatting troubles. A single round of behavioral analysis may omit consideration of long-term user behavior, and reflection on this is something the Reflector can do.
@@ 4.4. Conversational Recommendation (CR)
Conversational recommendation systems (CRS) engage users in a dialogue to refine their preferences and deliver more accurate suggestions.
In conversational scenarios, the user’s input text is not necessarily explicitly instructive. Hence, the Task Interpreter can help translate the conversation history into a more concise and clear task prompt. In addition, the user’s input requirements may contain information unknown to the Manager, e.g., a product that has not sold on the platform. In this case, the Searcher can help the Manager understand what the user mentioned.
Beyond the abovementioned applications, our framework can support other scenarios by customizing the prompts and agents’ collaboration modes.
@@ 5. Interface Demonstration
The interface can be divided into two main panels: 1) The configuration panel, where users can select different tasks to tackle, such as ”Conversational Recommendation.” Users can also customize different systems and configuration files for the task execution. 2) The interaction panel, where the whole collaboration process takes place. Agents with different abilities would complete this task collaboratively.
In Figure 2, the user has expressed a preference for the movie ”Schindler’s List” and seeks recommendations for similar historical movies. As shown in subfigure (a), the Interpreter summarizes this input and translates it into a clearer task successfully. Then, the Manager calls for the help of the Searcher for two rounds, searching for movies about history and movies similar to ”Schindler’s List”. According to all the information, the Manager gives the final recommendation movie ”Amistad”.
@@ 6. Conclusion
In this work, we propose a novel LLM-based multi-agent collaboration framework for recommendation, called MACRec. Unlike existing studies on using agents for user/item simulation, we directly tackle recommendation tasks through the collaborative efforts of various specialized agents. We present applications of MACRec on four different recommendation tasks, including rating prediction, sequential recommendation, conversational recommendation, and explanation generation of recommendation results. Moreover, We developed an online web interface for MACRec, visualizing how agents collaboratively tackle tasks.

### v3

# MACRec: a Multi-Agent Collaboration Framework for Recommendation

@@ Abstract.
Multi-agents; Large Language Models; Recommender Systems
journalyear: 2024copyright: rightsretainedconference: Proceedings of the 47th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 14–18, 2024; Washington, DC, USAbooktitle: Proceedings of the 47th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’24), July 14–18, 2024, Washington, DC, USAdoi: 10.1145/3626772.3657669isbn: 979-8-4007-0431-4/24/07ccs: Information systems Recommender systems
@@ 1. Introduction
Researchers have started to utilize the capabilities of agents to solve recommendation tasks. Existing work like  (Zhang et al., 2023c; Wang et al., 2023b; Zhang et al., 2023b) primarily focuses on employing agents for simulating user or item behaviors, providing insights into user preferences but falling short of integration into RSs. On the other hand, some studies (Wang et al., 2023a; Huang et al., 2023) attempt to leverage the capabilities of agents to directly build a recommender, primarily using one single agent with planning and memory components and auxiliary tools (e.g., search engine). However, there are various complex decision-making tasks in recommendation scenarios (Sun et al., 2023, 2024), on which single-agent instances are unable to perform well. Multi-agent collaboration, which is near to human workflows, is believed to accomplish complex tasks better with collective intelligence. Although work (Shu et al., 2023) proposes a multi-agent recommendation framework, it only has limited agent types and a fixed collaboration mode.
@@ 3.1. Framework Overview
The following sections will detail each agent’s specific characteristics and functions. [^1] Code is available at https://github.com/wzf2000/MACRec.
@@ 3.2.1. Manager
For the given task, the Manager would assign sub-tasks to other agents and complete the main execution process. It oversees the collaboration among all other agents.
@@ 3.2.2. Reflector
The Reflector will step in when the Manager is about to perform the second or more runs on the same task input. If the Reflector judges that the answer given by the Manager has no room for improvement, the Manager will no longer perform the current run. Otherwise, the Reflector will further summarize where the Manager can be improved, e.g., not considering the few highly rated items/movies in the user’s historical interactions.
@@ 3.2.5. Task Interpreter
The Task Interpreter will get the conversation history when starts running. Since conversation histories can be long, the Task Interpreter will only get the last part of the history. The Task Interpreter also has access to call the text summarization tool to get a more concise overview of the history. Eventually, the Task Interpreter will give a specific description of the task requirements that will be used to guide the subsequent runs of the Manager.
@@ 4.2. Sequential Recommendation (SR)
Modeling of user’s long-term and short-term interests is important in sequential recommendation tasks. Hence, the User Analyst’s role is self-evident. The number of relevant items in the sequence is significantly higher than the rating prediction task. It is hard to ask the Item Analyst to analyze every item that appeared in either the history or the candidate set. Moreover, given that the answers to the sequential recommendation task are much more complex (i.e., a ranking order of the candidate set), the Reflector can help to avoid the Manager getting into formatting troubles. A single round of behavioral analysis may omit consideration of long-term user behavior, and reflection on this is something the Reflector can do.
@@ 4.4. Conversational Recommendation (CR)
Conversational recommender systems engage users in a dialogue to refine their preferences and deliver more accurate suggestions.
In conversational scenarios, the user’s input text is not necessarily explicitly instructive. Hence, the Task Interpreter can help translate the conversation history into a more concise and clear task prompt. In addition, the user’s input requirements may contain information unknown to the Manager. In this case, the Searcher can help the Manager understand what the user mentioned.
Beyond the abovementioned applications, our framework can support other scenarios by customizing the configurations.
@@ 5. Interface Demonstration
The interface can be divided into two main panels. 1) Configuration panel, where users can select different tasks to tackle, such as ”Rating Prediction.” Users can also customize different systems and configuration files for the task execution. 2) Interaction panel, where the whole collaboration process takes place. Agents with different abilities would complete the task collaboratively.
In Figure 2, the user has expressed a preference for the movie ”Schindler’s List” and seeks recommendations for similar historical movies. The Interpreter summarizes this input and translates it into a clearer task. Then, the Manager calls for the help of the Searcher for two rounds, searching for movies about history and movies similar to ”Schindler’s List”. According to all the information, the Manager gives the final recommendation movie ”Amistad”.
@@ 6. Conclusion
In this work, we propose a novel LLM-based multi-agent collaboration framework for recommendation, called MACRec. Unlike existing studies on using agents for user/item simulation, we directly tackle recommendation tasks through the collaboration of various agents. We present applications of MACRec on four different recommendation tasks. Moreover, We developed an online web interface for MACRec, visualizing how agents work collaboratively.